### v1

# VideoRFT: Incentivizing Video Reasoning Capability in MLLMs via Reinforced Fine-Tuning

Abstract
Reinforcement fine-tuning (RFT) has shown great promise in achieving human-level reasoning capabilities of Large Language Models (LLMs), and has recently been extended to MLLMs. Nevertheless, reasoning about videos, which is a fundamental aspect of human intelligence, remains a persistent challenge due to the complex logic, temporal and causal structures inherent in video data. To fill this gap, we propose VideoRFT, a novel approach that extends the RFT paradigm to cultivate human-like video reasoning capabilities in MLLMs. VideoRFT follows the standard two-stage scheme in RFT: supervised fine-tuning (SFT) with chain-of-thought (CoT) annotations, followed by reinforcement learning (RL) to improve generalization. A central challenge to achieve this in the video domain lies in the scarcity of large-scale, high-quality video CoT datasets. We address this by building a fully automatic CoT curation pipeline. First, we devise a cognition-inspired prompting strategy to elicit a reasoning LLM to generate preliminary CoTs based solely on rich, structured, and literal representations of video content. Subsequently, these CoTs are revised by a visual-language model conditioned on the actual video, ensuring visual consistency and reducing visual hallucinations. This pipeline results in two new datasets – VideoRFT-CoT-102K for SFT and VideoRFT-RL-310K for RL. To further strength the RL phase, we introduce a novel semantic-consistency reward that explicitly promotes the alignment between textual reasoning with visual evidence. This reward encourages the model to produce coherent, context-aware reasoning outputs grounded in visual input. Extensive experiments show that VideoRFT achieves state-of-the-art performance on six video reasoning benchmarks.
1 Introduction
The ability to reason about complex videos lies at the core of human cognitive development spelke2007core . Humans, even infants, exhibit a remarkable capacity to understand videos – recognizing what has happened, inferring what will happen next, and explaining why events occur. Replicating this capability in AI systems has become a central goal in video understanding, and has been extensively studied in the field of computer vision over the part decade chen2024mecd ; min2024morevqa ; yi2019clevrer ; TRN . Despite the progress, most AI models remain limited to perceptual-level understanding and struggle to reason about video content with the depth, efficiency, and interpretability characteristic of human cognition.
Figure 1: Overview of VideoRFT. (a) An example of CoT derived from VideoRFT. (b) VideoRFT achieves leading performance in six datasets. (c) The two-stage RFT underpins the training of VideoRFT.
Recently, the rapid development of video MLLMs, such as Video-ChatGPT VideoChatGPT , VILA VILA , Video-LLaVA VideoLLaVA , have significantly advanced the capabilities of video understanding. However, these models are predominantly answer-driven, i.e., producing answers without explicitly revealing the reasoning process. VoT VoT overcomes this by introducing a human-like reasoning framework that structures the task of video reasoning into five predefined steps using fixed templates. Nevertheless, such a rigid, template-based approach stands in contrast to the flexibility of human cognition, which is able to adapt the reasoning process based on perceptual inputs GroundedCognition ; DualProcessCognition .
In contrast, the very recent advancements, e.g., OpenAI-o1 o1 , DeepSeek-R1 DeepseekR1 , and Kimi-1.5 Kimi15 , have shifted focus towards building LLMs that think before answering. These models show strong proficiency in interpreting complex problems, performing multi-step reasoning, and ultimately arriving at correct answers. A key enabler to such capabilities is reinforcement fine-tuning (RFT) luong2024reft , which typically commences with a warm-up phase of supervised fine-tuning using CoTs, and subsequently refines the model through reinforcement learning algorithms (e.g., PPO PPO , GRPO shao2024deepseekmath ). Beyond the language domain, pioneering efforts have extended RFT to MLLMs to enhance image-based capabilities VisionR1 ; VisualRFT ; ReasonRFT ; R1OneVision ; R1VL ; R1Omni , and some works VideoR1 ; VideoChatR1 ; TinyLLaVA-Video-R1 that concurrently with ours show the potential of RFT in the video domain. However, there is a critical challenge remain unsolved: current video CoT datasets lack the complexity and granularity necessary for advanced video reasoning, which fundamentally limits the ability of models to emulate human-level cognitive capabilities. Moreover, how to ensure that reasoning outputs are faithfully grounded in visual evidence remains underexplored in these works.
Motivated by above analysis, we propose VideoRFT, a novel reinforcement fine-tuning framework to incentive the video reasoning capability in MLLMs (see Fig. 1). To overcome the scarcity of video CoTs, we develop a scalable, cognitively inspired pipeline for high-quality video CoT dataset construction. Specifically, we first extract structured textual descriptions from videos using a MLLM to capture fine-grained visual details. The descriptions are then fed into a reasoning-capable LLM (e.g., DeepSeek-R1), which generates initial CoTs through blind reasoning—relying solely on textual descriptions. However, due to the lack of visual context, these CoTs often suffer from inconsistencies and hallucinations Faithful ; HallucinationSurvey . To resolve this, we develop a cross-modal revision stage, wherein a MLLM refines the initial CoTs by taking original video content, ensuring consistency with visual evidence. Based on this pipeline, we construct two large-scale datasets, i.e., VideoRFT-CoT-102K and VideoRFT-RL-310K, which together support the RFT process in VideoRFT.
Furthermore, to strengthen the RL phase, we develop a novel semantic-consistency reward that explicitly enhances the visual faithfulness of reasoning output in MLLMs. Our key observation is that the reasoning traces of MLLMs are typically structured into three consecutive parts for question parsing, video describing, and abstract reasoning. While the question parsing and abstract reasoning components are not necessarily grounded in the visual input, the video describing part should closely align with the actual visual semantics. Based on this insight, our semantic-consistency reward measures the alignment between the token representations of the video description part and the visual features of the input video. This reward is integrated into the GRPO algorithm to guide MLLMs towards generating visually grounded outputs.
Contributions of this work. We propose VideoRFT, a novel framework that extends RFT to MLLMs so as to emulate human-like video reasoning capabilities. To achieve this, we first establish a CoT foundation for video RFT by designing a cognitively inspired pipeline to curate large-scale, high-quality video CoT annotations. Furthermore, we introduce a novel semantic-consistency reward to explicitly guide the reasoning trajectories of MLLMs grounded in visual evidence, which enhances the effectiveness of RFT in cross-modal reasoning. Built on these contributions, VideoRFT favorably outperforms advanced competitors on a series of challenging video reasoning benchmarks.
2 VideoRFT CoT Dataset
We first present the construction of VideoRFT-COT and VideoRFT-RL to support RFT in MLLMs.
2.1 Data Collection
Figure 2: The distribution of data collection.
We extensively collect video question-answer data covering diverse modalities, task types, and cognitive skills. Given the scarcity of high-quality video data in certain domains (e.g., mathematics, science), we additionally incorporate carefully curated image-based instances. The final dataset contains 310K samples in total, supporting diverse answer formats, including multiple-choice (mc), numerical (num), free-form text generation (free), OCR (ocr), and regression (reg). As shown in Fig. 2, the samples are categorized into five groups according to the type of cognitive skills involving in the reasoning process:
  • General: Commonsense reasoning in open-domain temporal and causal contexts.
  • Mathematics: Supports symbolic reasoning and spatial alignment for multi-step logic tasks.
  • Science: Focuses on domain-specific reasoning in physics, chemistry, and medicine, emphasizing causal reasoning and conceptual abstraction.
  • Document: Targets structured visual parsing and information extraction from complex layouts.
  • Spatiotemporal: Involves motion prediction, spatial transformation, and relational reasoning.
2.2 Cognitively Inspired CoT Generation
To enable MLLMs to acquire human-like reasoning abilities, it is essential to construct a high-quality, cognitively grounded video CoT dataset. We propose an automated pipeline for generating such CoT data. As illustrated in Fig. 3, the pipeline comprises three major stages, and all the prompts used in the pipeline are provided in the supplementary material.
Structured Video Representation. For each video $v$, we generate semantically rich textual descriptions by prompting GPT-4o-mini gpt4o . The prompt $P_{\text{rep}}$ is carefully crafted to guide the model to (i) summarize video content with a high-level caption, and (ii) produce analytical, frame-level metadata for uniformly sampled video frames. Each frame is structured in a predefined JSON schema that includes timestamped captions and key visual elements such as objects, actions, scenes, spatial relations, and potential interactions. We denote the structured representation of $v$ as $S_{v}$.
Cognitively Inspired CoT Generation. Given the representation $S_{v}$ and a corresponding question $q$, we invoke a LLM, e.g., DeepSeek-R1 to answer the question and extracts its step-by-step reasoning outputs as the initial CoT, i.e., $\text{CoT}_{v}^{(0)}$:
$\small\text{CoT}_{v}^{(0)}=\text{LLM}(q,S_{v},P_{\text{cog}}).$ (1)
Here, $P_{\text{cog}}$ denotes a composite prompt consisting of five sub-prompts i.e., $P_{\text{cog}}=[p_{s},p_{t},p_{a},p_{v},p_{r}]$, each guiding a distinct stage of the reasoning trajectory in a manner that mimics human cognitive processing. Concretely, they are defined as follows. (1) Simulated observation prompt ($p_{s}$): Instruct the model to simulate viewing the entire video to form an initial high-level understanding of the video scene. (2) Task understanding prompt ($p_{t}$): Encourage analysis of the question $q$ to infer the task type (e.g., fact, reason, causal relationship). (3) Selective focus prompt ($p_{a}$): Direct attention to specific temporal segments of the video relevant to $q$. (4) Visual reasoning prompt ($p_{v}$): Ground the reasoning process in visual content, encouraging analysis over objects, actions, spatial-temporal relations, and event transitions. (5) Reflective answering prompt ($p_{r}$): Guide the model to derive the final answer, optionally incorporating self-verification or reflection to ensure reasoning quality.
Cross-modal CoT Refinement. A key limitation in the initial CoTs is that it might suffer from visual hallucinations due to the lack of visual cues in Eq. 1. To resolve this issue, we introduce a cross-modal refinement strategy to revise the CoT so that it aligns better with the actual video input. Specifically, we prompt a MLLM, i.e., Qwen2.5-VL Qwen25VL , to compare the initial CoT with the video $v$, identify inconsistencies, and perform necessary revisions:
$\small\text{CoT}_{v}=\text{MLLM}(v,\text{CoT}_{v}^{(0)},{P}_{\text{cross}}).$ (2)
Here the prompt ${P}_{\text{cross}}$ is designed to guide the MLLM: i) verify the cross-modal alignment of $\text{CoT}_{v}^{(0)}$ with the content of video $v$, ii) localize and explain any visual-textual inconsistencies, and iii) revise the CoT to enhance visual grounding while preserving its original logical structure.
Finally, we apply a filtering stage to ensure the factual correctness of the resulting CoT annotations. For structured tasks with clear ground-truth labels, we directly discard samples with incorrect final answers. For open-ended tasks, we remove samples exhibiting low semantic consistency (measured by CLIP CLIP ) between the generated answer and the reference answer. This filtering process ensures that the resulting CoT dataset maintains high quality and factual reliability. After filtering, 102K high-confidence samples are retained from the initial pool of 310K, forming VideoRFT-CoT-102K for supervised fine-tuning.
[…]
Figure 3: Illustration of the pipeline for cognitively inspired CoT generation.
2.3 Data Analysis
(a) Distribution of CoT token length in VideoRFT-CoT-102K
(b) Word cloud in VideoRFT-CoT-102K
(c) Distribution of CoT token length in Video-R1
(d) Word cloud in Video-R1
Figure 4: Comparison of CoT dataset in VideoRFT-CoT-102K and Video-R1.
Fig. 4 presents a comparative analysis of the CoTs in our proposed VideoRFT-CoT-102K and Video-R1 VideoR1 . As shown in Fig. 4 (a), the CoTs in our dataset exhibit a broader and longer token length distribution compared to those in Video-R1 (Fig. 4(c)), indicating that our VideoRFT-CoT-102K contains more elaborate, fine-grained, and nuanced reasoning processes. Additionally, the word cloud in Fig. 4 (b) reveal that CoTs in VideoRFT-CoT-102K are dominated by dynamic, video-centric concepts such as “video”, “main”, “happen”, and “first”. The lexical profile reflects an emphasis on narrative structure and temporal progression, which are the key characteristics of complex video understanding. In contrast, Video-R1 (Fig. 4 (d)) features frequent references to static or declarative content, e.g., “diagram”, “image”, “plant”, and “Earth”, suggesting a stronger bias toward factual descriptions rather than deep reasoning. These results highlight that VideoRFT offers greater expressiveness in reasoning depth, and aligns more closely with the demands of real-world video reasoning tasks. Hence, it provides a better foundation for training video MLLMs with advanced reasoning capabilities.
Figure 5: Illustrations of (a) rule-based RL, and (b) the computation of semantic-consistency reward $R_{s}$. The reasoning outputs are color-coded to highlight question parsing (green), video description (red) and abstract reasoning (blue). Only the red part is involved in the computation of $R_{s}$ (see §3.2).
3 Video Reinforcement Fine-Tuning
This section presents our approach for video reinforcement fine-tuning as shown in Fig. 5. We first make a brief overview of Group Relative Policy Optimization (GRPO) shao2024deepseekmath in §3.1, and then elaborate on the proposed rule-based reward for efficient reinforcement fine-tuning in §3.2.
3.1 Group Relative Policy Optimization
GRPO shao2024deepseekmath is a computationally efficient rule-based RL algorithm designed specifically for training large reasoning models. Unlike traditional RL methods such as PPO PPO which require four models (policy, value, reward, and reference), GRPO simplifies the approach by eliminating the value model, significantly reducing memory requirements and training complexity. GRPO operates by generating $K$ multiple candidate responses $\{o_{1},o_{2},\ldots,o_{K}\}$ for each query $q$. These responses are then evaluated using defined reward functions, yielding rewards $\{r_{1},r_{2},\ldots,r_{K}\}$. Afterwards, these rewards are normalized to calculate the advantage $A_{i}$ for each response as:
$\small A_{i}=\frac{r_{i}-{\texttt{mean}(\{r_{1},r_{2},\cdots,r_{K}\})}}{{\texttt{std}(\{r_{1},r_{2},\cdots,r_{K}\})}},$ (3)
where mean and std denote the mean and standard deviation of the rewards, respectively. Subsequently, the model is optimized through maximization of the following objective:
$\small\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{[q,\{o_{i}\}]}\frac{1}{K}\sum_{i=1}^{K}\left[\min\left(\frac{\pi_{\theta}}{\pi_{\theta_{\text{old}}}}A_{i},\text{clip}\left(\frac{\pi_{\theta}}{\pi_{\theta_{\text{old}}}},1\!-\!\epsilon,1\!+\!\epsilon\right)A_{i}\right)-\beta\mathbb{D}_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})\right],$ (4)
where $\theta$ denotes model parameters to be updated, $\pi_{\theta}$ and $\pi_{\theta_{\text{old}}}$ are the current and old policy, $\pi_{\text{ref}}$ indicates the reference policy, and $\beta$ is the KL divergence regularization coefficient. The $\epsilon$ is a regularization coefficient, preventing the policy from deviating too far from the reference model.
3.2 Rule-based Reward Modeling in VideoRFT
The rewards in Eq. 3 are derived from rule-based reward functions, which represent a foundational step in rule-based RL by simply evaluating whether model predictions exactly match ground-truth answers. Two highly resilient rule-based rewards are the Format Reward and Accuracy Reward, which are consistently utilized in DeepSeek-R1 and its follow-ups. However, in the context of cross-modal reasoning, these rewards are insufficient to provide explicit guidance to MLLMs towards visually grounded reasoning. To address this limitation, we introduce a semantic-consistency reward, which enforces the grounding of generated reasoning content in the visual input.
3.2.1 Semantic-Consistency Reward
The reward is motivated by the observation that the reasoning trace generated by video MLLMs typically consists of three distinct parts, i.e., question parsing, video describing and abstract reasoning, as shown in Fig. 5. Among them, the video describing stage the model’s understanding of visual content, which is the foundation for subsequent reasoning. Therefore, the reward is designed to selectively promote alignment between this stage and the input video.
Formally, to isolate the video describing sentence from the generated response, we apply a regular expression to locate the first full stop. Empirically, the portion following this punctuation corresponds to the model’s visual interpretation. From this point, we extract a fixed-length span of $M$ tokens $t$, and encode it using SigLIP SIGLIP ’s text encoder: $\bm{t}\!=\!\text{SigLIP}_{\text{text}}(t)$. Additionally, we uniformly sample $F$ frames $\{v^{(0)},\cdots,v^{(F)}\}$ from video $v$, and compute the visual representation of each frame $v^{(i)}$ via SigLIP’s image encoder: $\bm{v}^{(i)}\!=\!\text{SigLIP}_{\text{image}}(v^{(i)})$. Then the final video representation ${\bm{v}}$ is naturally obtained by averaging the frame embeddings: ${\bm{v}}\!=\!\frac{1}{F}\sum_{i=1}^{F}\bm{v}^{(i)}$. We then define the semantic-consistency reward as:
$\small R_{s}=\min\left(1,w\times\max\left(\text{cos}\left(\bm{t},{\bm{v}}\right),0\right)\right),$ (5)
where $\text{cos}(\cdot,\cdot)$ denotes cosine similarity, and $w\!=\!2$ is a scaling constant. The $\max(\cdot,0)$ ensures non-negativity of the reward, while the $\min(\cdot,1)$ stabilizes training by bounding the reward. This stage-aware formulation allows us to reward only the part of reasoning tied to visual comprehension, without penalizing abstract reasoning that appropriately extends beyond the visual scope. The result is enhanced semantic fidelity, reduced hallucinations, and improved alignment during RL.
3.3 Overall Reward
VideoRFT uses three types of rewards for RL:
  • Format Reward. During RL, we incorporate the widely-used format reward to guide the model in generating its reasoning process and final answer in a structured format. This reward, denoted as $R_{f}$, ensures that the model’s output adhere to a predefined structure: reasoning process must be enclosed within <think></think> tags, and answers within <answer></answer> tags. Compliance is verified via regular expression matching, and a binary reward is assigned accordingly.
  • Accuracy Reward. To provide reliable supervision across heterogeneous tasks, we adopt task-specific accuracy metrics: exact match for multiple-choice and numerical questions, ROUGE for open-ended generation, Word Error Rate (WER) for OCR tasks, and a scaled relative accuracy for regression problems. These tailored evaluations ensure the reward $R_{a}$ aligns with each task.
  • Semantic-Consistency Reward. The semantic-consistency reward $R_{s}$, defined in Eq. 5, promotes alignment between the reasoning text and the input visual information.
The overall reward $R$ for a sample is computed as follows:
$\small R=R_{f}+R_{a}+\mathbb{1}[R_{a}>0]\cdot R_{s},$ (6)
where $\mathbb{1}[R_{a}>0]$ is the indicator function that returns 1 if $R_{a}\!>\!0$ and 0 otherwise. It serves as a gate to control that the semantic-consistency reward $R_{s}$ is only activated when the accuracy reward $R_{a}$ is non-zero, which avoids reinforcing semantically plausible but factually incorrect reasoning.
4 Experiment
4.1 Experimental Setup
Benchmark and Metric. Following previous works VideoR1 ; LLaVAOneVision ; TinyLLaVA-Video-R1 , we evaluate our approach on six video reasoning and understanding benchmarks: VSI-Bench VSIBench , VideoMMMU VideoMMMU , MMVU MMVU , MVBench MVBench , TempCompass TempCompass and VideoMME VideoMME , covering spatial reasoning, knowledge-intensive video QA, temporal logic, and general video understanding. Following conventions, we only use the subset of multiple-choice samples in MMVU, and VideoMME is evaluated without subtitles. Average accuracy is adopted as the evaluation metric.
Model Training. We follow the RFT to train VideoRFT in two stages: the warm-up SFT stage and the rule-based RL stage. Specifically, the SFT stage equips the model with the ability to generate correct responses for diverse questions, and is trained based VideoRFT-CoT-102K. The rule-based RL stage is based on VideoRFT-RL-310K using the reward in Eq. 6 to optimize structured reasoning and ensure factual validity.
Implementation Details. We use Qwen2.5-VL-7B Qwen25VL as the base model and train VideoRFT on 8 NVIDIA A800 GPUs, with 80GB per each. For efficiency, the video input is limited to 16 frames of resolution $128\times 28\times 28$ during training, where $28\!\times\!28$ is the size of each image patch, and $128$ denotes the number of patches. During inference, we increase the number of frames to $32$ and the resolution to $256\times 28\times 28$. For efficiency, we use a lightweight version of SigCLIP with 400M parameters in computing the semantic-consistency reward. The entire model is trained for one epoch of SFT followed by 1K steps of RL.
Table 1: Performance Comparison. The best results are highlighted in bold. : Results are obtained using larger input resolutions, up to $768\!\times\!28\!\times\!28$ and 768 sampled frames, while ours are $256\!\times\!28\!\times\!28$ and $32$.
  Video Reasoning Video Understanding
Model Pub VSI. VideoMMMU MMVU MV. TempCom. VideoMME
$\bullet$ Proprietary Models
GPT-4o gpt4o 34.0 61.2 75.4 - - 71.9
$\bullet$ Open-Sourced Models
LLaMA-VID LLaMA-VID ECCV 24 - - - 41.9 45.6 -
ShareGPT4Video ShareGPT4Video NeurIPS 24 - - - 51.2 - 39.9
VideoLLaMA2 VideoLLaMA2 arXiv 24.06 - - 44.8 54.6 - 47.9
LongVA-7B LongVA arXiv 24.06 29.2 23.9 - - 56.9 52.6
VILA-1.5-8B VILA CVPR 24 28.9 20.8 - - 58.8 -
LLaVA-OneVision-7B LLaVAOneVision arXiv 24.08 32.4 33.8 49.2 56.7 - 58.2
mPLUG-Owl3-8B mPLUGOwl3 ICLR 25 - - - 54.5 - 53.5
Qwen2.5-VL-7B Qwen25VL arXiv 25.02 31.8 47.4 61.3 59.4 69.2 52.8
$\bullet$ Concurrent R1-based Models
Video-R1 VideoR1 arXiv 25.03 35.8 52.3 63.8 63.9 73.2 59.3
TinyLLaVA-Video-R1 TinyLLaVA-Video-R1 arXiv 25.04 - - 46.9 - 49.5 46.6
VideoChat-R1 VideoChatR1 arXiv 25.04 - - - 67.9 - -
VideoRFT 36.8 51.1 68.5 62.1 73.7 59.8
4.2 Main Result
As in Table 1, we compare VideoRFT against a variety of baselines, including proprietary models (i.e., GPT-4o gpt4o ), open-sourced MLLMs (e.g., Qwen2.5-VL Qwen25VL , VILA VILA , LongVA LongVA ), and contemporaneous models (e.g., Video-R1 VideoR1 , TinyLLaVA-Video-R1 TinyLLaVA-Video-R1 , VideoChat-R1 VideoChatR1 ).
Several key observations can be drawn from the results. First, compared to our base model, i.e., Qwen2.5-VL-7B, VideoRFT achieves significant improvements across all six benchmarks, e.g., +5.0% on VSI-Bench, +7.2% on MMVU, and +7.0 on VideoMME. This demonstrates the effectiveness of our approach in incentivizing video reasoning capabilities in MLLMs. Moreover, VideoRFT consistently outperforms all non-RL open-sourced MLLMs. Second, VideoRFT surpasses the proprietary GPT-4o on VSI-Bench by +2.8%, highlighting the strong potential of RFT in in bridging the performance gap with closed-source models in the task of video reasoning. Third, when compared to contemporaneous works, our model delivers the best overall performance, ranking first on four out of six benchmarks. This validates the superiority and generalization ability of our method in comparison with recent endeavors.
4.3 Diagnostic Experiment
To gain deeper insights into VideoRFT, we conduct a set of diagnostic experiments, as in Table 2.
Training Paradigm. To validate the effectiveness of the RFT training paradigm in our approach, we build two baseline approaches: i.e., SFT only and RL only. The former trains the model solely with supervised fine-tuning on VideoRFT-CoT-102K, while the latter, also known as the “zero” model in DeepSeek-R1, relies exclusively on RL without prior SFT. As seen from Table 2, the RL only surpasses the SFT only counterpart on four out of six datasets, indicating the capability of RL in stimulating more generalized reasoning capability. When combining both stages as in RFT, our full model VideoRFT achieves the best results, substantially outperforming the two baselines across all datasets. This highlights the complementary strengths of SFT for stable initialization and RL for reasoning enhancement in tackling video reasoning.
Reward Modeling. We further examine the effect of the reward defined in Eq. 6. Specifically, we compare two ablated variants: the first uses only uses the Format Reward and Accuracy Reward ($R\!=\!R_{f}\!+\!R_{a}$), while the second incorporates all three rewards directly ($R\!=\!R_{f}\!+\!R_{a}\!+\!R_{f}$). As seen, adding the semantic-consistency reward $R_{s}$ consistently improves performance over the first variant, validating its effectiveness. Finally, our full reward formulation, which conditionally activates $R_{s}$ via gating, achieves the best overall results. Notably, it brings substantial gains especially for video reasoning benchmarks, i.e., +2.2% on VSI-Bench, +0.9% on VideoMMMU and +3.3% on MMVU.
Training Curve. Fig. 6 shows the training dynamics of VideoRFT, including accuracy reward, response length, and semantic-consistency reward. The accuracy reward in Fig. 6 (a) shows an overall increasing trend, which means that GRPO contributes to gradually enhancing the model’s accuracy. Moreover, we observe steady improvements in response length and semantic reward. In the early stages, both metrics increase significantly, indicating that the model begins generating more complex and semantically grounded reasoning chains.
Table 2: Diagnostic experiments for VideoRFT.
  Video Reasoning Video Understanding
Model VSI-Bench VideoMMMU MMVU MVBench TempCompass VideoMME
$\bullet$ Training Paradigm
SFT only 31.7 48.5 60.5 57.0 68.4 54.1
RL only 32.1 47.4 63.5 59.2 70.8 51.9
$\bullet$ Reward Modeling
$R=R_{f}\!+\!R_{a}$ 33.2 49.1 66.4 61.1 72.4 58.5
$R=R_{f}\!+\!R_{a}\!+\!R_{s}$ 34.6 50.2 65.2 61.4 73.9 56.3
$\bullet$ Full Model
VideoRFT 36.8 51.1 68.5 62.1 73.7 59.8
(a) Accuracy reward.
(b) Response length.
(c) Semantic-consistency reward.
Figure 6: Training curves of VideoRFT in the RL phase.
Figure 7: Illustration of reasoning traces derived from VideoRFT in VSI-Bench.
Aha Moment in VideoRFT. Fig. 7 exhibits an Aha Moment in VideoRFT, where it behaves in a human-like manner by pausing to double-check its inference before finalizing the answer, as seen in the phrase “Wait, let me double-check it”. Such behavior suggests that the model is not simply recalling learned patterns, but is instead engaging in internal feedback loops to re-evaluate evidence and refine its inference.
5 Related Work
5.1 Multimodal Reasoning in MLLMs
Enabling reasoning in MLLMs has become a central objective in recent research li2024enhancing ; AoTD ; MMCoT ; DDCoT . In the image domain, early works such as MMCoT MMCoT and DDCoT DDCoT disentangle perception and reasoning by treating visual understanding as input prompts for subsequent inference. In the video domain, VoT VoT and STEP STEP decompose video reasoning into predefined stages, employing template-based prompting to facilitate multi-step inference. DoraemonGPT DoraemonGPT models video understanding through symbolic memory and external tool sequences, yet still follows a modular reasoning paradigm. While these methods offer structured supervision, their rigid designs often limit generalization across diverse temporal and causal scenarios. Recently, rule-based RL have emerged as a promising paradigm for promoting multimodal reasoning in MLLMs. Pioneering efforts such as Visual-RFT VisualRFT , R1-VL R1VL , and Reason-RFT ReasonRFT directly adapt rule-based RL to image perception tasks. Follow-ups like Vision-R1 VisionR1 and R1-OneVision R1OneVision further demonstrate its effectiveness in enabling CoT reasoning on images. Concurrently, this paradigm has also been explored for video understanding VideoChatR1 ; VideoR1 ; TinyLLaVA-Video-R1 . Despite encouraging progress, these methods face a fundamental bottleneck: the lack of large-scale, high-quality video CoT datasets, which limits the full potential of RFT in the video domain. Our work addresses this gap by proposing a scalable and cognitively inspired pipeline to automatically mine high-quality CoT annotations for videos. Beyond this, we introduce a novel reward modeling strategy based on cross-modal semantic consistency, which explicitly guides MLLMs to generate visually grounded reasoning traces, which proves to be highly effective in improving model performance.
5.2 Multimodal CoT Dataset Construction
CoT has proven effective for enhancing the reasoning capabilities of LLMs by encouraging step-by-step reasoning CoT ; AutomaticCoT . Constructing high-quality CoT data in multimodal settings, particularly for video reasoning, remains a major challenge due to the temporal complexity and visual ambiguity of video data VideoEspresso ; VideoCoT ; VisualCoT . Recent works have explored CoT construction in both image and video domains. LLaVA-CoT llavacot , Vision-R1 VisionR1 , and R1-OneVision R1OneVision simply convert visual inputs to textual descriptions before reasoning. This often leads to hallucinations and weak semantic alignment. Video-R1 VideoR1 adopts a simplistic prompting strategy that encourages MLLMs to generate CoT by inserting “let me think", “wait". into responses. However, such CoTs merely mimic the surface form of human thinking without engaging in genuine reasoning. VideoEspresso VideoEspresso generates CoT data by prompting GPT-4o with a small set of selected key frames. Due to the sparse visual context and reliance on a text-only model, the generated CoTs often lack grounding in the actual video content and are prone to hallucinations. In contrast, our CoT data combines the reasoning ability of reasoning LLMs and the multimodal ability of MLLMs, ensuring the reasoning depth and visual grounding of CoT data. Moreover, we use cognition-inspired prompts to enable the reasoning model to generate CoT data that is more in line with human cognition.
6 Conclusion
In this work, we introduce VideoRFT, a novel approach for incentivizing cognitive video reasoning capabilities in MLLMs through reinforced fine-tuning. To accomplish this, we propose a cross-modal pipeline that generates high-quality cognitive video CoT data simulating human reasoning processes, resulting in two large-scale, high-quality datasets: VideoRFT-CoT-102K and VideoRFT-RL-310K. Furthermore, to strengthen the RL phase, we develop semantic-consistency guided reward modeling to explicitly encourage the alignment between reasoning traces and visual evidence. Extensive experiments across six video reasoning benchmarks demonstrate that VideoRFT consistently surpasses a variety of advanced MLLMs. We expect this work lay a foundation for future efforts in RFT-based video reasoning.

### v3

@@
 A central challenge to achieve this in the video domain lies in the scarcity of large-scale, high-quality video CoT datasets.
-We address this by building a fully automatic CoT curation pipeline.
+We address this by building a multi-expert, cognition-inspired CoT curation pipeline.
 First, we devise a cognition-inspired prompting strategy to elicit a reasoning LLM to generate preliminary CoTs based solely on rich, structured, and literal representations of video content.
-Subsequently, these CoTs are revised by a visual-language model conditioned on the actual video, ensuring visual consistency and reducing visual hallucinations.
-This pipeline results in two new datasets – VideoRFT-CoT-102K for SFT and VideoRFT-RL-310K for RL.
-To further strength the RL phase, we introduce a novel semantic-consistency reward that explicitly promotes the alignment between textual reasoning with visual evidence.
+Subsequently, these CoTs are revised by a MLLM conditioned on the actual video, ensuring visual consistency and reducing visual hallucinations.
+This pipeline results in two new datasets, i.e.VideoRFT-CoT-102K for SFT and VideoRFT-RL-310K for RL.
+To further strengthen the RL phase, we introduce a novel semantic-consistency reward that explicitly promotes the alignment between textual reasoning and visual evidence.
 This reward encourages the model to produce coherent, context-aware reasoning outputs grounded in visual input.
@@
 Humans, even infants, exhibit a remarkable capacity to understand videos – recognizing what has happened, inferring what will happen next, and explaining why events occur.
-Replicating this capability in AI systems has become a central goal in video understanding, and has been extensively studied in the field of computer vision over the part decade chen2024mecd ; min2024morevqa ; yi2019clevrer ; TRN .
-Despite the progress, most AI models remain limited to perceptual-level understanding and struggle to reason about video content with the depth, efficiency, and interpretability characteristic of human cognition.
+Replicating this capability in AI systems has become a central goal in video understanding, and has been extensively studied in the field of computer vision over the past decade chen2024mecd ; min2024morevqa ; yi2019clevrer ; TRN .
+Despite the progress, most AI models remain limited to perceptual-level understanding and struggle to reason about video content with the depth, efficiency, and interpretability that characteristic of human cognition.
 Figure 1: Overview of VideoRFT. (a) An example of CoT derived from VideoRFT. (b) VideoRFT achieves leading performance in six datasets. (c) The two-stage RFT underpins the training of VideoRFT.
-Recently, the rapid development of video MLLMs, such as Video-ChatGPT VideoChatGPT , VILA VILA , Video-LLaVA VideoLLaVA , have significantly advanced the capabilities of video understanding.
+Recently, the rapid development of video MLLMs, such as Video-ChatGPT VideoChatGPT , VILA VILA , Video-LLaVA VideoLLaVA , has significantly advanced the capabilities of video understanding.
 However, these models are predominantly answer-driven, i.e., producing answers without explicitly revealing the reasoning process.
@@
 A key enabler to such capabilities is reinforcement fine-tuning (RFT) luong2024reft , which typically commences with a warm-up phase of supervised fine-tuning using CoTs, and subsequently refines the model through reinforcement learning algorithms (e.g., PPO PPO , GRPO shao2024deepseekmath ).
-Beyond the language domain, pioneering efforts have extended RFT to MLLMs to enhance image-based capabilities VisionR1 ; VisualRFT ; ReasonRFT ; R1OneVision ; R1VL ; R1Omni , and some works VideoR1 ; VideoChatR1 ; TinyLLaVA-Video-R1 that concurrently with ours show the potential of RFT in the video domain.
-However, there is a critical challenge remain unsolved: current video CoT datasets lack the complexity and granularity necessary for advanced video reasoning, which fundamentally limits the ability of models to emulate human-level cognitive capabilities.
+Beyond the language domain, pioneering efforts have extended RFT to MLLMs to enhance image-based capabilities VisionR1 ; VisualRFT ; ReasonRFT ; R1OneVision ; R1VL ; R1Omni , and some works VideoR1 ; VideoChatR1 ; TinyLLaVA-Video-R1 tha, concurrently with ours, show the potential of RFT in the video domain.
+However, there is a critical challenge remaining unsolved: current video CoT datasets lack the complexity and granularity necessary for advanced video reasoning, which fundamentally limits the ability of models to emulate human-level cognitive capabilities.
 Moreover, how to ensure that reasoning outputs are faithfully grounded in visual evidence remains underexplored in these works.
-Motivated by above analysis, we propose VideoRFT, a novel reinforcement fine-tuning framework to incentive the video reasoning capability in MLLMs (see Fig.
+Motivated by the above analysis, we propose VideoRFT, a novel reinforcement fine-tuning framework to incentivize the video reasoning capability in MLLMs (see Fig.
 1).
-To overcome the scarcity of video CoTs, we develop a scalable, cognitively inspired pipeline for high-quality video CoT dataset construction.
-Specifically, we first extract structured textual descriptions from videos using a MLLM to capture fine-grained visual details.
-The descriptions are then fed into a reasoning-capable LLM (e.g., DeepSeek-R1), which generates initial CoTs through blind reasoning—relying solely on textual descriptions.
-However, due to the lack of visual context, these CoTs often suffer from inconsistencies and hallucinations Faithful ; HallucinationSurvey .
-To resolve this, we develop a cross-modal revision stage, wherein a MLLM refines the initial CoTs by taking original video content, ensuring consistency with visual evidence.
+To overcome the scarcity of video CoTs, we develop a scalable, cognitively inspired pipeline that integrates multiple expert models to collaboratively construct high-quality video CoT datasets.
+Specifically, we first employ a specialized MLLM to extract structured textual descriptions from videos, capturing fine-grained visual details.
+These descriptions are then processed by a reasoning-capable LLM (e.g., DeepSeek-R1), which generates initial CoTs through blind reasoning—relying solely on textual input.
+However, due to the lack of direct visual grounding, such CoTs often contain inconsistencies and hallucinations (HallucinationSurvey, ).
+To mitigate this issue, we introduce a cross-modal revision stage, wherein a MLLM refines the initial CoTs by incorporating the original video, ensuring consistency with visual evidence.
 Based on this pipeline, we construct two large-scale datasets, i.e., VideoRFT-CoT-102K and VideoRFT-RL-310K, which together support the RFT process in VideoRFT.
-Furthermore, to strengthen the RL phase, we develop a novel semantic-consistency reward that explicitly enhances the visual faithfulness of reasoning output in MLLMs.
+Furthermore, to strengthen the RL phase, we develop a novel semantic-consistency reward that explicitly enhances the visual faithfulness of reasoning outputs in MLLMs.
 Our key observation is that the reasoning traces of MLLMs are typically structured into three consecutive parts for question parsing, video describing, and abstract reasoning.
@@
 Based on this insight, our semantic-consistency reward measures the alignment between the token representations of the video description part and the visual features of the input video.
-This reward is integrated into the GRPO algorithm to guide MLLMs towards generating visually grounded outputs.
+This reward is integrated into the GRPO algorithm to guide MLLMs toward generating visually grounded outputs.
 Contributions of this work.
@@
 As shown in Fig.
-2, the samples are categorized into five groups according to the type of cognitive skills involving in the reasoning process:
+2, the samples are categorized into five groups according to the type of cognitive skills involved in the reasoning process:
   • General: Commonsense reasoning in open-domain temporal and causal contexts.
@@
 Cognitively Inspired CoT Generation.
-Given the representation $S_{v}$ and a corresponding question $q$, we invoke a LLM, e.g., DeepSeek-R1 to answer the question and extracts its step-by-step reasoning outputs as the initial CoT, i.e., $\text{CoT}_{v}^{(0)}$:
+Given the representation $S_{v}$ and a corresponding question $q$, we invoke a LLM, e.g., DeepSeek-R1 to answer the question and extract its step-by-step reasoning outputs as the initial CoT, i.e., $\text{CoT}_{v}^{(0)}$:
 $\small\text{CoT}_{v}^{(0)}=\text{LLM}(q,S_{v},P_{\text{cog}}).$ (1)
@@
 Concretely, they are defined as follows.
-(1) Simulated observation prompt ($p_{s}$): Instruct the model to simulate viewing the entire video to form an initial high-level understanding of the video scene.
+(1) Simulated observation prompt ($p_{s}$): Instruct the model to simulate viewing the entire video to form an initial high-level understanding.
 (2) Task understanding prompt ($p_{t}$): Encourage analysis of the question $q$ to infer the task type (e.g., fact, reason, causal relationship).
@@
 Cross-modal CoT Refinement.
-A key limitation in the initial CoTs is that it might suffer from visual hallucinations due to the lack of visual cues in Eq.
+A key limitation in the initial CoTs is that they might suffer from visual hallucinations due to the lack of visual cues in Eq.
 1.
@@
 For structured tasks with clear ground-truth labels, we directly discard samples with incorrect final answers.
-For open-ended tasks, we remove samples exhibiting low semantic consistency (measured by CLIP CLIP ) between the generated answer and the reference answer.
+For open-ended tasks, we remove samples exhibiting low semantic consistency (measured by CLIP CLIP ) between the generated and the reference answer.
 This filtering process ensures that the resulting CoT dataset maintains high quality and factual reliability.
 After filtering, 102K high-confidence samples are retained from the initial pool of 310K, forming VideoRFT-CoT-102K for supervised fine-tuning.
-[…]
 Figure 3: Illustration of the pipeline for cognitively inspired CoT generation.
@@
 Additionally, the word cloud in Fig.
-4 (b) reveal that CoTs in VideoRFT-CoT-102K are dominated by dynamic, video-centric concepts such as “video”, “main”, “happen”, and “first”.
+4 (b) reveals that CoTs in VideoRFT-CoT-102K are dominated by dynamic, video-centric concepts such as “video”, “main”, “happen”, and “first”.
 The lexical profile reflects an emphasis on narrative structure and temporal progression, which are the key characteristics of complex video understanding.
@@
 4 (d)) features frequent references to static or declarative content, e.g., “diagram”, “image”, “plant”, and “Earth”, suggesting a stronger bias toward factual descriptions rather than deep reasoning.
-These results highlight that VideoRFT offers greater expressiveness in reasoning depth, and aligns more closely with the demands of real-world video reasoning tasks.
+These results highlight that VideoRFT offers greater expressiveness in reasoning depth and aligns more closely with the demands of real-world video reasoning tasks.
 Hence, it provides a better foundation for training video MLLMs with advanced reasoning capabilities.
@@
 5.
-We first make a brief overview of Group Relative Policy Optimization (GRPO) shao2024deepseekmath in §3.1, and then elaborate on the proposed rule-based reward for efficient reinforcement fine-tuning in §3.2.
+We first provide a brief overview of Group Relative Policy Optimization (GRPO) shao2024deepseekmath in §3.1, and then elaborate on the proposed rule-based reward for efficient reinforcement fine-tuning in §3.2.
 3.1 Group Relative Policy Optimization
 GRPO shao2024deepseekmath is a computationally efficient rule-based RL algorithm designed specifically for training large reasoning models.
-Unlike traditional RL methods such as PPO PPO which require four models (policy, value, reward, and reference), GRPO simplifies the approach by eliminating the value model, significantly reducing memory requirements and training complexity.
-GRPO operates by generating $K$ multiple candidate responses $\{o_{1},o_{2},\ldots,o_{K}\}$ for each query $q$.
+Unlike traditional RL methods such as PPO PPO , which require four models (policy, value, reward, and reference), GRPO simplifies the approach by eliminating the value model, significantly reducing memory requirements and training complexity.
+GRPO operates by generating $K$ candidate responses $\{o_{1},o_{2},\ldots,o_{K}\}$ for each query $q$.
 These responses are then evaluated using defined reward functions, yielding rewards $\{r_{1},r_{2},\ldots,r_{K}\}$.
@@
 $\small\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{[q,\{o_{i}\}]}\frac{1}{K}\sum_{i=1}^{K}\left[\min\left(\frac{\pi_{\theta}}{\pi_{\theta_{\text{old}}}}A_{i},\text{clip}\left(\frac{\pi_{\theta}}{\pi_{\theta_{\text{old}}}},1\!-\!\epsilon,1\!+\!\epsilon\right)A_{i}\right)-\beta\mathbb{D}_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})\right],$ (4)
-where $\theta$ denotes model parameters to be updated, $\pi_{\theta}$ and $\pi_{\theta_{\text{old}}}$ are the current and old policy, $\pi_{\text{ref}}$ indicates the reference policy, and $\beta$ is the KL divergence regularization coefficient.
-The $\epsilon$ is a regularization coefficient, preventing the policy from deviating too far from the reference model.
+where $\theta$ denotes model parameters to be updated, $\pi_{\theta}$ and $\pi_{\theta_{\text{old}}}$ are the current and old policy model, $\pi_{\text{ref}}$ indicates the reference policy, and $\beta$ is the KL divergence regularization coefficient.
+Here, $\epsilon$ is a regularization coefficient, preventing the policy from deviating too far from the reference model.
 3.2 Rule-based Reward Modeling in VideoRFT
@@
 5.
-Among them, the video describing stage the model’s understanding of visual content, which is the foundation for subsequent reasoning.
+Among them, the video describing stage represents the model’s understanding of visual content, which is the foundation for subsequent reasoning.
 Therefore, the reward is designed to selectively promote alignment between this stage and the input video.
@@
 Empirically, the portion following this punctuation corresponds to the model’s visual interpretation.
-From this point, we extract a fixed-length span of $M$ tokens $t$, and encode it using SigLIP SIGLIP ’s text encoder: $\bm{t}\!=\!\text{SigLIP}_{\text{text}}(t)$.
+From this point, we extract a fixed-length span of $M$ tokens, denoted $t$, and encode it using SigLIP SIGLIP ’s text encoder: $\bm{t}\!=\!\text{SigLIP}_{\text{text}}(t)$.
 Additionally, we uniformly sample $F$ frames $\{v^{(0)},\cdots,v^{(F)}\}$ from video $v$, and compute the visual representation of each frame $v^{(i)}$ via SigLIP’s image encoder: $\bm{v}^{(i)}\!=\!\text{SigLIP}_{\text{image}}(v^{(i)})$.
-Then the final video representation ${\bm{v}}$ is naturally obtained by averaging the frame embeddings: ${\bm{v}}\!=\!\frac{1}{F}\sum_{i=1}^{F}\bm{v}^{(i)}$.
+Then the final video representation $\bm{v}$} is naturally obtained by averaging the frame embeddings: ${\bm{v}}\!=\!\frac{1}{F}\sum_{i=0}^{F}\bm{v}^{(i)}$.
 We then define the semantic-consistency reward as:
 $\small R_{s}=\min\left(1,w\times\max\left(\text{cos}\left(\bm{t},{\bm{v}}\right),0\right)\right),$ (5)
-where $\text{cos}(\cdot,\cdot)$ denotes cosine similarity, and $w\!=\!2$ is a scaling constant.
+where $\text{cos}(\cdot,\cdot)$ denotes cosine similarity, and w=2w\!=\!2 is a scaling constant.
 The $\max(\cdot,0)$ ensures non-negativity of the reward, while the $\min(\cdot,1)$ stabilizes training by bounding the reward.
@@
 During RL, we incorporate the widely-used format reward to guide the model in generating its reasoning process and final answer in a structured format.
-This reward, denoted as $R_{f}$, ensures that the model’s output adhere to a predefined structure: reasoning process must be enclosed within <think></think> tags, and answers within <answer></answer> tags.
+This reward, denoted as $R_{f}$, ensures that the model’s output adheres to a predefined structure: the reasoning process must be enclosed within <think></think> tags, and answers within <answer></answer> tags.
 Compliance is verified via regular expression matching, and a binary reward is assigned accordingly.
@@
 where $\mathbb{1}[R_{a}>0]$ is the indicator function that returns 1 if $R_{a}\!>\!0$ and 0 otherwise.
-It serves as a gate to control that the semantic-consistency reward $R_{s}$ is only activated when the accuracy reward $R_{a}$ is non-zero, which avoids reinforcing semantically plausible but factually incorrect reasoning.
+It serves as a gate to control whether the semantic-consistency reward $R_{s}$ is only activated when the accuracy reward $R_{a}$ is non-zero, which avoids reinforcing semantically plausible but factually incorrect reasoning.
 4 Experiment
@@
 We follow the RFT to train VideoRFT in two stages: the warm-up SFT stage and the rule-based RL stage.
-Specifically, the SFT stage equips the model with the ability to generate correct responses for diverse questions, and is trained based VideoRFT-CoT-102K.
+Specifically, the SFT stage equips the model with the ability to generate correct responses for diverse questions, and is trained based on VideoRFT-CoT-102K.
 The rule-based RL stage is based on VideoRFT-RL-310K using the reward in Eq.
 6 to optimize structured reasoning and ensure factual validity.
+The RL training is implemented using the HuggingFace TRL library trl , and our codebase is built upon Open-R1 openr1 .
 Implementation Details.
-We use Qwen2.5-VL-7B Qwen25VL as the base model and train VideoRFT on 8 NVIDIA A800 GPUs, with 80GB per each.
+We use Qwen2.5-VL-7B Qwen25VL as the base model and train VideoRFT on 8 NVIDIA A800 GPUs, with 80GB each.
 For efficiency, the video input is limited to 16 frames of resolution $128\times 28\times 28$ during training, where $28\!\times\!28$ is the size of each image patch, and $128$ denotes the number of patches.
@@
 The best results are highlighted in bold.
-: Results are obtained using larger input resolutions, up to $768\!\times\!28\!\times\!28$ and 768 sampled frames, while ours are $256\!\times\!28\!\times\!28$ and $32$.
+: Results are obtained using larger input resolutions, up to $768\!\times\!28\!\times\!28$ and 768 sampled frames, while ours are $256\!\times\!28\!\times\!28$ and 3232.
   Video Reasoning Video Understanding
-Model Pub VSI. VideoMMMU MMVU MV. TempCom. VideoMME
+Model Pub VSI. VideoMMMU MMVU MV. TempC. VideoMME
 $\bullet$ Proprietary Models
 GPT-4o gpt4o 34.0 61.2 75.4 - - 71.9
-$\bullet$ Open-Sourced Models
+$\bullet$ Open-Source Models
 LLaMA-VID LLaMA-VID ECCV 24 - - - 41.9 45.6 -
@@
 4.2 Main Result
-As in Table 1, we compare VideoRFT against a variety of baselines, including proprietary models (i.e., GPT-4o gpt4o ), open-sourced MLLMs (e.g., Qwen2.5-VL Qwen25VL , VILA VILA , LongVA LongVA ), and contemporaneous models (e.g., Video-R1 VideoR1 , TinyLLaVA-Video-R1 TinyLLaVA-Video-R1 , VideoChat-R1 VideoChatR1 ).
+As in Table 1, we compare VideoRFT against a variety of baselines, including proprietary models (i.e., GPT-4o gpt4o ), Open-Source MLLMs (e.g., Qwen2.5-VL Qwen25VL , VILA VILA , LongVA LongVA ), and contemporaneous models (e.g., Video-R1 VideoR1 , TinyLLaVA-Video-R1 TinyLLaVA-Video-R1 , VideoChat-R1 VideoChatR1 ).
 Several key observations can be drawn from the results.
-First, compared to our base model, i.e., Qwen2.5-VL-7B, VideoRFT achieves significant improvements across all six benchmarks, e.g., +5.0% on VSI-Bench, +7.2% on MMVU, and +7.0 on VideoMME.
+First, compared to our base model, i.e., Qwen2.5-VL-7B, VideoRFT achieves significant improvements across all six benchmarks, e.g., +5.0% on VSI-Bench, +7.2% on MMVU, and +7.0% on VideoMME.
 This demonstrates the effectiveness of our approach in incentivizing video reasoning capabilities in MLLMs.
-Moreover, VideoRFT consistently outperforms all non-RL open-sourced MLLMs.
-Second, VideoRFT surpasses the proprietary GPT-4o on VSI-Bench by +2.8%, highlighting the strong potential of RFT in in bridging the performance gap with closed-source models in the task of video reasoning.
+Moreover, VideoRFT consistently outperforms all non-RL Open-Source MLLMs.
+Second, VideoRFT surpasses the proprietary GPT-4o on VSI-Bench by +2.8%, highlighting the strong potential of RFT in bridging the performance gap with closed-source models in the task of video reasoning.
 Third, when compared to contemporaneous works, our model delivers the best overall performance, ranking first on four out of six benchmarks.
@@
 The former trains the model solely with supervised fine-tuning on VideoRFT-CoT-102K, while the latter, also known as the “zero” model in DeepSeek-R1, relies exclusively on RL without prior SFT.
-As seen from Table 2, the RL only surpasses the SFT only counterpart on four out of six datasets, indicating the capability of RL in stimulating more generalized reasoning capability.
+As seen from Table 2, the RL only surpasses the SFT only counterpart on four out of six datasets, indicating the capability of RL in stimulating more generalized reasoning capabilities.
 When combining both stages as in RFT, our full model VideoRFT achieves the best results, substantially outperforming the two baselines across all datasets.
@@
 6.
-Specifically, we compare two ablated variants: the first uses only uses the Format Reward and Accuracy Reward ($R\!=\!R_{f}\!+\!R_{a}$), while the second incorporates all three rewards directly ($R\!=\!R_{f}\!+\!R_{a}\!+\!R_{f}$).
+Specifically, we compare two ablated variants: the first uses only the Format Reward and Accuracy Reward ($R\!=\!R_{f}\!+\!R_{a}$), while the second incorporates all three rewards directly ($R\!=\!R_{f}\!+\!R_{a}\!+\!R_{s}$).
 As seen, adding the semantic-consistency reward $R_{s}$ consistently improves performance over the first variant, validating its effectiveness.
@@
 While these methods offer structured supervision, their rigid designs often limit generalization across diverse temporal and causal scenarios.
-Recently, rule-based RL have emerged as a promising paradigm for promoting multimodal reasoning in MLLMs.
+Recently, rule-based RL has emerged as a promising paradigm for promoting multimodal reasoning in MLLMs.
 Pioneering efforts such as Visual-RFT VisualRFT , R1-VL R1VL , and Reason-RFT ReasonRFT directly adapt rule-based RL to image perception tasks.
@@
 Our work addresses this gap by proposing a scalable and cognitively inspired pipeline to automatically mine high-quality CoT annotations for videos.
-Beyond this, we introduce a novel reward modeling strategy based on cross-modal semantic consistency, which explicitly guides MLLMs to generate visually grounded reasoning traces, which proves to be highly effective in improving model performance.
+Beyond this, we introduce a novel reward modeling strategy based on cross-modal semantic consistency, which explicitly guides MLLMs to generate visually grounded reasoning traces, and proves to be highly effective in improving model performance.
 5.2 Multimodal CoT Dataset Construction
@@
 This often leads to hallucinations and weak semantic alignment.
-Video-R1 VideoR1 adopts a simplistic prompting strategy that encourages MLLMs to generate CoT by inserting “let me think", “wait".
-into responses.
+Video-R1 VideoR1 adopts a simplistic prompting strategy that encourages MLLMs to generate CoT by inserting “let me think”, “wait”, etc.into responses.
 However, such CoTs merely mimic the surface form of human thinking without engaging in genuine reasoning.
@@
 Due to the sparse visual context and reliance on a text-only model, the generated CoTs often lack grounding in the actual video content and are prone to hallucinations.
-In contrast, our CoT data combines the reasoning ability of reasoning LLMs and the multimodal ability of MLLMs, ensuring the reasoning depth and visual grounding of CoT data.
+In contrast, our CoT data combines the reasoning abilities of reasoning LLMs and the multimodal abilities of MLLMs, ensuring the reasoning depth and visual grounding of CoT data.
 Moreover, we use cognition-inspired prompts to enable the reasoning model to generate CoT data that is more in line with human cognition.
@@
 Extensive experiments across six video reasoning benchmarks demonstrate that VideoRFT consistently surpasses a variety of advanced MLLMs.
-We expect this work lay a foundation for future efforts in RFT-based video reasoning.
+We expect this work to lay a foundation for future efforts in RFT-based video reasoning.
+Appendix A Additional Experiments
+A.1 Effect of Model Scaling on Reasoning Ability
+Table 3: Performance Comparison with Small-scale Models
+  Video Reasoning Video Understanding
+Model VSI. VideoMMMU MMVU MV. TempC. VideoMME
+TinyLLaVA-Video-R1-3B - - 46.9 - 49.5 46.6
+VideoRFT-3B 32.5 41.1 55.1 59.5 61.0 45.4
+VideoRFT-7B 36.8 51.1 68.5 62.1 73.7 59.8
+To further assess the scalability and robustness of our method under limited computational resources, we train an additional 3B variant of our model.
+While its performance is naturally lower than the original 7B version, the 3B model still performs competitively across benchmarks.
+Crucially, when compared to TinyLLaVA-Video-R1-3B, a contemporary 3B model specifically optimized for lightweight deployment, our 3B variant outperforms it, with gains of +8.2% on MMVU and +11.5% on TempCompass.
+This significant margin confirms the competitiveness of our approach even in small-scale settings, and underscores its strong generalization and reasoning abilities under resource-constrained conditions.
+A.2 Hyperparameter Analysis
+Table 4: Hyperparameter experiments for VideoRFT.
+  Video Reasoning Video Understanding
+Hyperparameter VSI-Bench VideoMMMU MMVU MVBench TempCompass VideoMME
+$w=1$ 34.2 49.2 67.9 62.6 73.5 61.1
+$w=2$ 36.8 51.1 68.5 62.1 73.7 59.8
+$w=3$ 35.6 50.1 67.9 60.8 72.3 60.4
+$w=4$ 35.4 49.7 67.6 62.0 73.1 59.1
+To evaluate the effect of the scaling factor $w$ in the semantic reward, we conducted a hyperparameter study across multiple benchmarks.
+As shown in Table˜4, the performance varies with different values of $w$, with w=2w{=}2 consistently achieving the best or near-best results across all benchmarks.
+This suggests that w=2w{=}2 provides a favorable balance between semantic alignment and factual precision.
+In comparison, both smaller (w=1w{=}1) and larger (w=3w{=}3 or w=4w{=}4) values lead to a slight decline in performance, implying that insufficient or excessive emphasis on semantic similarity may hinder the overall reasoning and understanding capabilities.
+These findings highlight that while performance is relatively stable across a range of scaling values, setting w=2w{=}2 consistently yields optimal results, demonstrating the effectiveness of moderate semantic weighting.
+Figure 8: Illustration of reasoning traces derived from VideoRFT in VideoMME.
+A.3 Aha Moment in VideoRFT
+Figure 8 illustrates additional instances of the “Aha Moment” behavior observed in VideoRFT.
+Prior to reaching a final decision, the model exhibits a human-like tendency to pause and scrutinize its intermediate reasoning steps.
+This reflective process, indicative of deliberative reasoning, is marked in red within the figure.
+Appendix B Limitations and Future Directions
+We consider this work a strong foundation for advancing video reasoning research using MLLMs. Several avenues warrant further investigation in future studies:
+  • Handling of Challenging Video Scenarios.
+While our model demonstrates strong performance across various benchmark tasks, its effectiveness may be affected under complex conditions such as rapid motion or severe visual occlusion.
+Incorporating finer-grained visual signals and higher frame-rate sampling may help mitigate these challenges, enabling more accurate and robust visual representations.
+  • CoT Data Reliance.
+The quality of the generated CoT annotations is closely tied to the capabilities of the underlying reasoning language model (e.g., DeepSeek-R1).
+Future work could explore leveraging more advanced reasoning models to further enhance the quality of CoT data, potentially leading to improved performance of VideoRFT.
+Appendix C Potential Social Impacts
+  • Positive Impacts.
+Enhanced video understanding enabled by VideoRFT can benefit various applications.
+In education, it facilitates the development of intelligent tutoring systems using video content.
+In security, it improves the efficiency and accuracy of surveillance video analysis.
+Moreover, it supports content moderation by aiding in the detection and filtering of inappropriate material.
+  • Negative Impacts.
+There exists a risk of misuse.
+Misinterpretation of video content due to over-reliance on automated analysis could lead to the spread of misinformation.
+Additionally, in surveillance scenarios, the deployment of such systems may raise concerns regarding privacy and ethical use.
+Appendix D Detailed Prompt Used in CoT Generation
+Here, we provide detailed prompts for each step of the VideoRFT CoT generation process.
+D.1 Video CoT Generation
+[…]
+For all videos, we sample them at 1FPS and input them into GPT-4o-mini, and generate a structured representation according to the prompt $P_{\text{rep}}$:
+[…]
+You are a video analysis assistant designed to produce rich, analytical per-frame captions from video inputs.
+[…]
+<Input Video>
+[Uncaptioned image] </Input Video>
+Task: 1. Overall Video Caption: video_caption: A concise 1-2 sentence (20-30 words) summary capturing the main theme or action of the video. 2. Per-Frame Metadata:
+Uniformly sample frames and wrap them into a JSON list. Each element must include:
+timestamp: "$H$:$M$:SS", aligned with the sampling interval.
+caption: 2-3 sentences (30-50 words) describing the scene in detail.
+key_elements: an object with fields: objects: list of detected objects/entities (strings) actions: list of ongoing actions or movements (strings) scene: e.g.“kitchen”, “urban street” notable_features: list of distinctive colors, textures, or patterns (strings) spatial_relations: list of spatial relationships (e.g.“cup on table”, “person left of car”) human_attributes: object or null.
+If present, include: * gender: “male”, “female”, “unknown” * clothing: brief description * posture: “standing”, “sitting” potential_interactions: list of possible interactions (strings)
+General Instructions: Temporal Consistency: Reference continuing actions from the previous frame and highlight any changes.
+Uncertainty: If confidence<<0.6 or object visibility<<50%, append [Uncertain] to the caption.
+Implied Actions: Describe preparatory movements (e.g.“hand reaching toward door handle” vs.
+“holding door handle”).
+Output Requirements: Wrap all per-frame objects into a single JSON list matching the number of sampled frames.
+[…]
+We invoke DeepSeek-R1 to answer the question and extracts its step-by-step reasoning outputs with the prompt $P_{\text{cog}}$ as the initial CoT.
+[…]
+You are an AI assistant helping a user answer questions about a video.
+When the user asks a question, you respond by imagining you are watching the video with full attention, just like a human would.
+Your task is to reason visually and logically about the video content to answer the user’s question.
+[…]
+Video content: <Overall Video Caption>
+Frame-level metadata: <Per-Frame Metadata> Question: <Question> Please think about this question as if you were a human pondering deeply.
+It’s encouraged to include self-reflection or verification in the reasoning process.
+<Corresponding Answer Format Template>
+The <Corresponding Answer Format Template> in the prompt is dynamically selected from the following templates based on the question type:
+Table 5: Answer Format Templates for Different Question Types
+Question Type Template
+Multiple Choice Please provide only the single option letter (e.g., A, B, C, D, etc.) within the <answer> </answer> tags.
+Numerical Please provide the numerical value within the <answer> </answer> tags.
+OCR Please transcribe text from the image/video clearly and provide your text answer within the <answer> </answer> tags.
+Free-form Please provide your text answer within the <answer> </answer> tags.
+Regression Please provide the numerical value within the <answer> </answer> tags.
+[…]
+We employ a cross-modal refinement process to ensure the CoT aligns with the video content, using the prompt $P_{\text{cross}}$:
+[…]
+You are a multimodal reasoning expert.
+Your task is to revise hallucinations and errors in the chain-of-thought (CoT) based on the visual content of the provided video.
+Do not significantly alter the original CoT logic or content, and ensure the final conclusion remains the same.
+[…]
+<Input Video>
+[Uncaptioned image] </Input Video>
+Question: <Question> Original CoT: <Original CoT> Output format:
+Strictly follow this format. Return only the revised CoT and no additional explanation:
+<think>[Revised CoT]</think>
+D.2 Image CoT Generation
+In addition to the main video data in the VideoRFT-CoT-102K, we have also designed specific prompts specifically for image data, and used the similarly CoT generation process to generate CoT.
+[…]
+Due to the differences among different image datasets (the task focuses of the datasets are different), we designed different prompts for each image dataset in structured image representation phase.
+Common System Prompt for All Image Representations You are a vision-language expert. Your task is to analyze the provided image and output a detailed, modular description in JSON format.
+The following prompts are used for different image datasets, all sharing the above system prompt:
+[…]
+<Input Image>
+[Uncaptioned image] </Input Image>
+You are given a natural photograph.
+Output only valid JSON with two keys: 1.
+Overall Image Caption: a 2-3 sentence narrative describing the scene, objects, actions, and context.
+2.
+Image Metadata: an object containing: objects: list of objects with id, type, color, size, and bbox text: list of text elements with content and bbox scene_context: environment and activity relations: list of subject-predicate-object relations
+[…]
+<Input Image>
+[Uncaptioned image] </Input Image>
+You are given a synthetic 3D scene.
+Output only valid JSON with: 1.
+Overall Image Caption: a brief summary of number of objects and overall layout.
+2.
+Image Metadata: objects: list of objects with id, shape, color, size, material, and coordinates spatial_relations: list of spatial relations between objects
+[…]
+<Input Image>
+[Uncaptioned image] </Input Image>
+You are given a line-drawing or diagram.
+Output only valid JSON with: 1.
+Overall Image Caption: a concise paragraph describing the diagram’s purpose.
+2.
+Image Metadata: primitives: list of geometric primitives with type, label, and bbox annotations: list of relations between primitives measurements: list of measurements with primitive IDs and values
+[…]
+<Input Image>
+[Uncaptioned image] </Input Image>
+You are given an image containing printed or handwritten text.
+Output only valid JSON with: 1.
+Overall Image Caption: one sentence summarizing the text context.
+2.
+Image Metadata: text_items: list of text elements with content, bbox, and style
+[…]
+<Input Image>
+[Uncaptioned image] </Input Image>
+You are given a multi-panel scientific figure.
+Output only valid JSON with: 1.
+Overall Image Caption: a paragraph overviewing the figure’s subject.
+2.
+Image Metadata: panels: list of panels with elements and process arrows
+[…]
+<Input Image>
+[Uncaptioned image] </Input Image>
+You are given a chart image.
+Output only valid JSON with: 1.
+Overall Image Caption: a 1-2 sentence summary of chart type and key trend.
+2.
+Image Metadata: chart_type: type of chart axes: list of axis information series: list of data series legend: list of legend entries
+[…]
+<Input Image>
+[Uncaptioned image] </Input Image>
+You are given a textbook problem image.
+Output only valid JSON with: 1.
+Overall Image Caption: a summary of the problem context.
+2.
+Image Metadata: equations: list of equations with LaTeX and bbox diagram_parts: list of diagram elements givens: list of given values
+[…]
+<Input Image>
+[Uncaptioned image] </Input Image>
+You are given an indoor scene or floorplan image.
+Output only valid JSON with: 1.
+Overall Image Caption: a 2-3 sentence narrative of the space.
+2.
+Image Metadata: rooms_or_sections: list of room information furniture: list of furniture items with position and orientation annotations: list of structural elements
+[…]
+We invoke DeepSeek-R1 to answer the question and extracts its step-by-step reasoning outputs with the prompt $P_{\text{cog}}$ as the initial CoT.
+[…]
+You are an AI assistant helping a user answer questions about an image.
+When the user asks a question, you respond by imagining you are looking at the image with full attention, just like a human would.
+Your task is to reason visually and logically about the image content to answer the user’s question.
+[…]
+Image content: <Overall Image Caption>
+Image metadata: <Image Metadata> Question: <Question> Please think about this question as if you were a human pondering deeply.
+It’s encouraged to include self-reflection or verification in the reasoning process.
+<Corresponding Answer Format Template>
+[…]
+We employ a cross-modal refinement process to ensure the CoT aligns with the image content, using the prompt $P_{\text{cross}}$:
+[…]
+You are a multimodal reasoning expert.
+Your task is to revise hallucinations and errors in the chain-of-thought (CoT) based on the provided image.
+Do not significantly alter the original CoT logic or content, and ensure the final conclusion remains the same.
+[…]
+<Input Image>
+[Uncaptioned image] </Input Image>
+Question: <Question> Original CoT: <Original CoT> Output format:
+Strictly follow this format. Return only the revised CoT and no additional explanation:
+<think>[Revised CoT]</think>

### v4

@@
 A central challenge to achieve this in the video domain lies in the scarcity of large-scale, high-quality video CoT datasets.
-We address this by building a multi-expert, cognition-inspired CoT curation pipeline.
+We address this by building a multi-expert-driven, cognition-inspired CoT curation pipeline.
 First, we devise a cognition-inspired prompting strategy to elicit a reasoning LLM to generate preliminary CoTs based solely on rich, structured, and literal representations of video content.
@@
 Humans, even infants, exhibit a remarkable capacity to understand videos – recognizing what has happened, inferring what will happen next, and explaining why events occur.
-Replicating this capability in AI systems has become a central goal in video understanding, and has been extensively studied in the field of computer vision over the past decade chen2024mecd ; min2024morevqa ; yi2019clevrer ; TRN .
-Despite the progress, most AI models remain limited to perceptual-level understanding and struggle to reason about video content with the depth, efficiency, and interpretability that characteristic of human cognition.
+Replicating this capability in AI systems has become a central goal in video understanding, and has been extensively studied in the field of computer vision over the past decade chen2024mecd ; min2024morevqa ; yi2020clevrer ; TRN .
+Despite the progress, most AI models remain limited to perceptual-level understanding and struggle to reason about video content with the depth, efficiency, and interpretability that are characteristic of human cognition.
 Figure 1: Overview of VideoRFT. (a) An example of CoT derived from VideoRFT. (b) VideoRFT achieves leading performance in six datasets. (c) The two-stage RFT underpins the training of VideoRFT.
 Recently, the rapid development of video MLLMs, such as Video-ChatGPT VideoChatGPT , VILA VILA , Video-LLaVA VideoLLaVA , has significantly advanced the capabilities of video understanding.
-However, these models are predominantly answer-driven, i.e., producing answers without explicitly revealing the reasoning process.
+However, these models are predominantly answer-driven, i.e., they produce answers without explicitly revealing the reasoning process.
 VoT VoT overcomes this by introducing a human-like reasoning framework that structures the task of video reasoning into five predefined steps using fixed templates.
-Nevertheless, such a rigid, template-based approach stands in contrast to the flexibility of human cognition, which is able to adapt the reasoning process based on perceptual inputs GroundedCognition ; DualProcessCognition .
+Nevertheless, such a rigid, template-based approach stands in contrast to the flexibility of human cognition, which enables adaptive reasoning processes based on perceptual inputs GroundedCognition ; DualProcessCognition .
 In contrast, the very recent advancements, e.g., OpenAI-o1 o1 , DeepSeek-R1 DeepseekR1 , and Kimi-1.5 Kimi15 , have shifted focus towards building LLMs that think before answering.
@@
 A key enabler to such capabilities is reinforcement fine-tuning (RFT) luong2024reft , which typically commences with a warm-up phase of supervised fine-tuning using CoTs, and subsequently refines the model through reinforcement learning algorithms (e.g., PPO PPO , GRPO shao2024deepseekmath ).
-Beyond the language domain, pioneering efforts have extended RFT to MLLMs to enhance image-based capabilities VisionR1 ; VisualRFT ; ReasonRFT ; R1OneVision ; R1VL ; R1Omni , and some works VideoR1 ; VideoChatR1 ; TinyLLaVA-Video-R1 tha, concurrently with ours, show the potential of RFT in the video domain.
+Beyond the language domain, pioneering efforts have extended RFT to MLLMs to enhance image-based capabilities VisionR1 ; VisualRFT ; ReasonRFT ; R1OneVision ; R1VL ; R1Omni , and some works VideoR1 ; VideoChatR1 ; TinyLLaVA-Video-R1 that concurrently with ours, show the potential of RFT in the video domain.
 However, there is a critical challenge remaining unsolved: current video CoT datasets lack the complexity and granularity necessary for advanced video reasoning, which fundamentally limits the ability of models to emulate human-level cognitive capabilities.
@@
 These descriptions are then processed by a reasoning-capable LLM (e.g., DeepSeek-R1), which generates initial CoTs through blind reasoning—relying solely on textual input.
-However, due to the lack of direct visual grounding, such CoTs often contain inconsistencies and hallucinations (HallucinationSurvey, ).
+However, due to the lack of direct visual grounding, such CoTs often contain inconsistencies and hallucinations HallucinationSurvey .
 To mitigate this issue, we introduce a cross-modal revision stage, wherein a MLLM refines the initial CoTs by incorporating the original video, ensuring consistency with visual evidence.
@@
 Furthermore, to strengthen the RL phase, we develop a novel semantic-consistency reward that explicitly enhances the visual faithfulness of reasoning outputs in MLLMs.
-Our key observation is that the reasoning traces of MLLMs are typically structured into three consecutive parts for question parsing, video describing, and abstract reasoning.
-While the question parsing and abstract reasoning components are not necessarily grounded in the visual input, the video describing part should closely align with the actual visual semantics.
+Our key observation is that the reasoning traces of MLLMs are typically structured into three consecutive parts: question parsing, video describing, and abstract reasoning.
+While the question parsing and abstract reasoning components are not necessarily grounded in the visual input, the video describing part should be closely aligned with the actual visual semantics.
 Based on this insight, our semantic-consistency reward measures the alignment between the token representations of the video description part and the visual features of the input video.
@@
 Given the scarcity of high-quality video data in certain domains (e.g., mathematics, science), we additionally incorporate carefully curated image-based instances.
-The final dataset contains 310K samples in total, supporting diverse answer formats, including multiple-choice (mc), numerical (num), free-form text generation (free), OCR (ocr), and regression (reg).
+The final dataset contains 310K samples in total, supporting diverse answer formats, including multiple-choice (mc), numerical (num), free-form text generation (free), optical character recognition (ocr), and regression (reg).
 As shown in Fig.
@@
   • General: Commonsense reasoning in open-domain temporal and causal contexts.
-  • Mathematics: Supports symbolic reasoning and spatial alignment for multi-step logic tasks.
-  • Science: Focuses on domain-specific reasoning in physics, chemistry, and medicine, emphasizing causal reasoning and conceptual abstraction.
+  • Mathematics: Symbolic reasoning and spatial alignment for multi-step logic tasks.
+  • Science: Domain-specific reasoning in physics, chemistry, and medicine, emphasizing causal reasoning and conceptual abstraction.
   • Document: Targets structured visual parsing and information extraction from complex layouts.
@@
 Here, $P_{\text{cog}}$ denotes a composite prompt consisting of five sub-prompts i.e., $P_{\text{cog}}=[p_{s},p_{t},p_{a},p_{v},p_{r}]$, each guiding a distinct stage of the reasoning trajectory in a manner that mimics human cognitive processing.
-Concretely, they are defined as follows.
-(1) Simulated observation prompt ($p_{s}$): Instruct the model to simulate viewing the entire video to form an initial high-level understanding.
+Concretely, these sub-prompts are defined as follows.
+(1) Simulated observation prompt ($p_{s}$): Instruct the model to simulate viewing the entire video and form an initial high-level understanding.
 (2) Task understanding prompt ($p_{t}$): Encourage analysis of the question $q$ to infer the task type (e.g., fact, reason, causal relationship).
@@
 $\small\text{CoT}_{v}=\text{MLLM}(v,\text{CoT}_{v}^{(0)},{P}_{\text{cross}}).$ (2)
-Here the prompt ${P}_{\text{cross}}$ is designed to guide the MLLM: i) verify the cross-modal alignment of $\text{CoT}_{v}^{(0)}$ with the content of video $v$, ii) localize and explain any visual-textual inconsistencies, and iii) revise the CoT to enhance visual grounding while preserving its original logical structure.
+Here the prompt ${P}_{\text{cross}}$ is designed to guide the MLLM to: (i) verify the cross-modal alignment of $\text{CoT}_{v}^{(0)}$ with the content of video $v$, (ii) localize and explain any visual-textual inconsistencies, and (iii) revise the CoT to enhance visual grounding while preserving its original logical structure.
 Finally, we apply a filtering stage to ensure the factual correctness of the resulting CoT annotations.
-For structured tasks with clear ground-truth labels, we directly discard samples with incorrect final answers.
-For open-ended tasks, we remove samples exhibiting low semantic consistency (measured by CLIP CLIP ) between the generated and the reference answer.
+For structured tasks with clear ground-truth labels, we directly exclude samples with incorrect final answers.
+For open-ended tasks, we remove samples exhibiting low semantic consistency (measured by CLIP CLIP ) between the generated answer and the reference answer.
 This filtering process ensures that the resulting CoT dataset maintains high quality and factual reliability.
@@
 2.3 Data Analysis
-(a) Distribution of CoT token length in VideoRFT-CoT-102K
+(a) Distribution of CoT tokens’ length in VideoRFT-CoT-102K
 (b) Word cloud in VideoRFT-CoT-102K
-(c) Distribution of CoT token length in Video-R1
+(c) Distribution of CoT tokens’ length in Video-R1
 (d) Word cloud in Video-R1
@@
 As shown in Fig.
-4 (a), the CoTs in our dataset exhibit a broader and longer token length distribution compared to those in Video-R1 (Fig.
+4 (a), the CoTs in our dataset exhibit a broader distribution and longer average token length compared to those in Video-R1 (Fig.
 4(c)), indicating that our VideoRFT-CoT-102K contains more elaborate, fine-grained, and nuanced reasoning processes.
@@
 4 (d)) features frequent references to static or declarative content, e.g., “diagram”, “image”, “plant”, and “Earth”, suggesting a stronger bias toward factual descriptions rather than deep reasoning.
-These results highlight that VideoRFT offers greater expressiveness in reasoning depth and aligns more closely with the demands of real-world video reasoning tasks.
+These results highlight that VideoRFT-CoT-102K offers greater expressiveness in reasoning depth and aligns more closely with the demands of real-world video reasoning tasks.
 Hence, it provides a better foundation for training video MLLMs with advanced reasoning capabilities.
@@
 Afterwards, these rewards are normalized to calculate the advantage $A_{i}$ for each response as:
-$\small A_{i}=\frac{r_{i}-{\texttt{mean}(\{r_{1},r_{2},\cdots,r_{K}\})}}{{\texttt{std}(\{r_{1},r_{2},\cdots,r_{K}\})}},$ (3)
+$\small A_{i}=\frac{r_{i}-{\texttt{mean}(\{r_{1},r_{2},\ldots,r_{K}\})}}{{\texttt{std}(\{r_{1},r_{2},\ldots,r_{K}\})}},$ (3)
 where mean and std denote the mean and standard deviation of the rewards, respectively. Subsequently, the model is optimized through maximization of the following objective:
 $\small\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{[q,\{o_{i}\}]}\frac{1}{K}\sum_{i=1}^{K}\left[\min\left(\frac{\pi_{\theta}}{\pi_{\theta_{\text{old}}}}A_{i},\text{clip}\left(\frac{\pi_{\theta}}{\pi_{\theta_{\text{old}}}},1-\!\epsilon,1+\epsilon\right)A_{i}\right)-\beta\mathbb{D}_{\text{KL}}(\pi_{\theta}||\pi_{\text{ref}})\right],$ (4)
-where $\theta$ denotes model parameters to be updated, $\pi_{\theta}$ and $\pi_{\theta_{\text{old}}}$ are the current and old policy model, $\pi_{\text{ref}}$ indicates the reference policy, and $\beta$ is the KL divergence regularization coefficient.
-Here, $\epsilon$ is a regularization coefficient, preventing the policy from deviating too far from the reference model.
+where $\theta$ denotes model parameters to be updated, $\pi_{\theta}$ and $\pi_{\theta_{\text{old}}}$ are the current and old policy model, $\pi_{\text{ref}}$ indicates the reference policy, $\beta$ is the KL divergence regularization coefficient, and $\epsilon$ is a regularization coefficient that prevents the policy from deviating too far from the reference model.
 3.2 Rule-based Reward Modeling in VideoRFT
@@
 Formally, to isolate the video describing sentence from the generated response, we apply a regular expression to locate the first full stop.
-Empirically, the portion following this punctuation corresponds to the model’s visual interpretation.
-From this point, we extract a fixed-length span of $M$ tokens, denoted $t$, and encode it using SigLIP SIGLIP ’s text encoder: $\bm{t}\!=\!\text{SigLIP}_{\text{text}}(t)$.
-Additionally, we uniformly sample $F$ frames $\{v^{(0)},\cdots,v^{(F)}\}$ from video $v$, and compute the visual representation of each frame $v^{(i)}$ via SigLIP’s image encoder: $\bm{v}^{(i)}\!=\!\text{SigLIP}_{\text{image}}(v^{(i)})$.
+Empirically, the text following this full stop corresponds to the model’s interpretation of visual content.
+From this point, we extract a fixed-length span of $M$ tokens, denoted $t_{[i,i+M]}$, and encode it using SigLIP SIGLIP ’s text encoder: $\bm{t}_{[i,i+M]}=\text{SigLIP}_{\text{text}}(t_{[i,i+M]})$.
+Additionally, we uniformly sample $F$ frames $\{v^{(0)},\ldots,v^{(F-1)}\}$ from video $v$, and compute the visual representation of each frame $v^{(i)}$ via SigLIP’s image encoder: $\bm{v}^{(i)}=\text{SigLIP}_{\text{image}}(v^{(i)})$.
 Then the final video representation $\bm{v}$} is naturally obtained by averaging the frame embeddings: ${\bm{v}}=\frac{1}{F}\sum_{i=0}^{F}\bm{v}^{(i)}$.
 We then define the semantic-consistency reward as:
-$\small R_{s}=\min\left(1,w\times\max\left(\text{cos}\left(\bm{t},{\bm{v}}\right),0\right)\right),$ (5)
-where $\text{cos}(\cdot,\cdot)$ denotes cosine similarity, and w=2w\!=\!2 is a scaling constant.
+$\small R_{s}=\min\left(1,w\times\max\left(\text{cos}\left(\bm{t}_{[i,i+M]},{\bm{v}}\right),0\right)\right),$ (5)
+where $\text{cos}(\cdot,\cdot)$ denotes cosine similarity, and $w=2$ is a scaling constant.
 The $\max(\cdot,0)$ ensures non-negativity of the reward, while the $\min(\cdot,1)$ stabilizes training by bounding the reward.
@@
   • Accuracy Reward.
-To provide reliable supervision across heterogeneous tasks, we adopt task-specific accuracy metrics: exact match for multiple-choice and numerical questions, ROUGE for open-ended generation, Word Error Rate (WER) for OCR tasks, and a scaled relative accuracy for regression problems.
+To provide reliable supervision across heterogeneous tasks, we adopt task-specific accuracy metrics: Exact Match for multiple-choice and numerical questions, ROUGE for open-ended generation, Word Error Rate for OCR tasks, and a scaled relative accuracy for regression problems.
 These tailored evaluations ensure the reward $R_{a}$ aligns with each task.
@@
 where $\mathbb{1}[R_{a}>0]$ is the indicator function that returns 1 if $R_{a}>0$ and 0 otherwise.
-It serves as a gate to control whether the semantic-consistency reward $R_{s}$ is only activated when the accuracy reward $R_{a}$ is non-zero, which avoids reinforcing semantically plausible but factually incorrect reasoning.
+This indicator function acts as a gate to ensure $R_{s}$ is activated only when $R_{a}$ is non-zero, thus avoiding the reinforcement of semantically plausible but factually incorrect reasoning.
 4 Experiment
@@
 Benchmark and Metric.
-Following previous works VideoR1 ; LLaVAOneVision ; TinyLLaVA-Video-R1 , we evaluate our approach on six video reasoning and understanding benchmarks: VSI-Bench VSIBench , VideoMMMU VideoMMMU , MMVU MMVU , MVBench MVBench , TempCompass TempCompass and VideoMME VideoMME , covering spatial reasoning, knowledge-intensive video QA, temporal logic, and general video understanding.
+Following previous works VideoR1 ; LLaVAOneVision ; TinyLLaVA-Video-R1 , we evaluate our approach on six video reasoning and understanding benchmarks: VSI-Bench VSIBench , VideoMMMU VideoMMMU , MMVU MMVU , MVBench MVBench , TempCompass TempCompass , and VideoMME VideoMME , covering spatial reasoning, knowledge-intensive video QA, temporal logic, and general video understanding.
 Following conventions, we only use the subset of multiple-choice samples in MMVU, and VideoMME is evaluated without subtitles.
@@
 We use Qwen2.5-VL-7B Qwen25VL as the base model and train VideoRFT on 8 NVIDIA A800 GPUs, with 80GB each.
-For efficiency, the video input is limited to 16 frames of resolution $128\times 28\times 28$ during training, where $28\!\times\!28$ is the size of each image patch, and $128$ denotes the number of patches.
+For efficiency, the video input is limited to 16 frames, with each frame processed into $128\times 28\times 28$ resolution during training, where $28\times 28$ is the size of each image patch, and $128$ denotes the number of patches.
 During inference, we increase the number of frames to $32$ and the resolution to $256\times 28\times 28$.
-For efficiency, we use a lightweight version of SigCLIP with 400M parameters in computing the semantic-consistency reward.
+For efficiency, we use a lightweight version of SigLIP with 400M parameters in computing the semantic-consistency reward.
 The entire model is trained for one epoch of SFT followed by 1K steps of RL.
@@
 VideoLLaMA2 VideoLLaMA2 arXiv 24.06 - - 44.8 54.6 - 47.9
-LongVA-7B LongVA arXiv 24.06 29.2 23.9 - - 56.9 52.6
+LongVA-7B LongVA TMLR 24 29.2 23.9 - - 56.9 52.6
 VILA-1.5-8B VILA CVPR 24 28.9 20.8 - - 58.8 -
-LLaVA-OneVision-7B LLaVAOneVision arXiv 24.08 32.4 33.8 49.2 56.7 - 58.2
+LLaVA-OneVision-7B LLaVAOneVision TMLR 24 32.4 33.8 49.2 56.7 - 58.2
 mPLUG-Owl3-8B mPLUGOwl3 ICLR 25 - - - 54.5 - 53.5
@@
 4.2 Main Result
-As in Table 1, we compare VideoRFT against a variety of baselines, including proprietary models (i.e., GPT-4o gpt4o ), Open-Source MLLMs (e.g., Qwen2.5-VL Qwen25VL , VILA VILA , LongVA LongVA ), and contemporaneous models (e.g., Video-R1 VideoR1 , TinyLLaVA-Video-R1 TinyLLaVA-Video-R1 , VideoChat-R1 VideoChatR1 ).
+As shown in Table 1, we compare VideoRFT against a variety of baselines, including proprietary models (i.e., GPT-4o gpt4o ), Open-Source MLLMs (e.g., Qwen2.5-VL Qwen25VL , VILA VILA , LongVA LongVA ), and contemporaneous models (e.g., Video-R1 VideoR1 , TinyLLaVA-Video-R1 TinyLLaVA-Video-R1 , VideoChat-R1 VideoChatR1 ).
 Several key observations can be drawn from the results.
@@
 To gain deeper insights into VideoRFT, we conduct a set of diagnostic experiments, as in Table 2.
+Training Data.
+To assess the impact of the cross-modal refinement in Eq.
+2 on data quality, we conduct SFT+RL using the data generated by Eq.
+1, denoted as $\text{CoT}_{v}^{(0)}$, resulting in a variant, i.e., without CoT Refinement.
+As shown in Table 2, this consistently leads to performance drops across all six benchmarks, i.e., -2.3% on VSI-Bench, -3.7% on MMVU, and -7.0% on VideoMME.
+This demonstrates that our cross-modal refinement effectively mitigates errors and hallucinations in the initial CoTs, thereby enhancing data quality and ultimately improving model performance.
 Training Paradigm.
@@
 Specifically, we compare two ablated variants: the first uses only the Format Reward and Accuracy Reward ($R=R_{f}+R_{a}$), while the second incorporates all three rewards directly ($R=R_{f}+R_{a}+R_{s}$).
-As seen, adding the semantic-consistency reward $R_{s}$ consistently improves performance over the first variant, validating its effectiveness.
-Finally, our full reward formulation, which conditionally activates $R_{s}$ via gating, achieves the best overall results.
-Notably, it brings substantial gains especially for video reasoning benchmarks, i.e., +2.2% on VSI-Bench, +0.9% on VideoMMMU and +3.3% on MMVU.
-Training Curve.
-Fig.
-6 shows the training dynamics of VideoRFT, including accuracy reward, response length, and semantic-consistency reward.
-The accuracy reward in Fig.
-6 (a) shows an overall increasing trend, which means that GRPO contributes to gradually enhancing the model’s accuracy.
-Moreover, we observe steady improvements in response length and semantic reward.
-In the early stages, both metrics increase significantly, indicating that the model begins generating more complex and semantically grounded reasoning chains.
+As seen from Table 2, adding the semantic-consistency reward $R_{s}$ consistently improves performance over the first variant, validating its effectiveness.
+Finally, our full reward formulation, which conditionally activates $R_{s}$ via a gating mechanism (i.e., the indicator function $\mathbb{1}[R_{a}>0]$), achieves the best overall results.
+Notably, it brings substantial gains especially for video reasoning benchmarks, i.e., +2.2% on VSI-Bench, +3.3% on MMVU, and +3.5% on VideoMME.
 Table 2: Diagnostic experiments for VideoRFT.
@@
 Model VSI-Bench VideoMMMU MMVU MVBench TempCompass VideoMME
+$\bullet$ Training Data
+w/o CoT Refinement 34.5 48.1 64.8 58.3 72.4 52.8
 $\bullet$ Training Paradigm
@@
 VideoRFT 36.8 51.1 68.5 62.1 73.7 59.8
-(a) Accuracy reward.
-(b) Response length.
-(c) Semantic-consistency reward.
-Figure 6: Training curves of VideoRFT in the RL phase.
 Figure 6: Illustration of reasoning traces derived from VideoRFT in VSI-Bench.
@@
 Follow-ups like Vision-R1 VisionR1 and R1-OneVision R1OneVision further demonstrate its effectiveness in enabling CoT reasoning on images.
-Concurrently, this paradigm has also been explored for video understanding VideoChatR1 ; VideoR1 ; TinyLLaVA-Video-R1 .
+Concurrently, this paradigm has also been explored for video understanding VideoR1 ; VideoChatR1 ; TinyLLaVA-Video-R1 .
 Despite encouraging progress, these methods face a fundamental bottleneck: the lack of large-scale, high-quality video CoT datasets, which limits the full potential of RFT in the video domain.
@@
 CoT has proven effective for enhancing the reasoning capabilities of LLMs by encouraging step-by-step reasoning CoT ; AutomaticCoT .
-Constructing high-quality CoT data in multimodal settings, particularly for video reasoning, remains a major challenge due to the temporal complexity and visual ambiguity of video data VideoEspresso ; VideoCoT ; VisualCoT .
+Constructing high-quality CoT data in multimodal settings, particularly for video reasoning, remains a major challenge due to the temporal complexity and visual ambiguity of video data VideoEspresso ; VisualCoT ; VideoCoT .
 Recent works have explored CoT construction in both image and video domains.
@@
 In this work, we introduce VideoRFT, a novel approach for incentivizing cognitive video reasoning capabilities in MLLMs through reinforced fine-tuning.
-To accomplish this, we propose a cross-modal pipeline that generates high-quality cognitive video CoT data simulating human reasoning processes, resulting in two large-scale, high-quality datasets: VideoRFT-CoT-102K and VideoRFT-RL-310K.
-Furthermore, to strengthen the RL phase, we develop semantic-consistency guided reward modeling to explicitly encourage the alignment between reasoning traces and visual evidence.
-Extensive experiments across six video reasoning benchmarks demonstrate that VideoRFT consistently surpasses a variety of advanced MLLMs.
+To accomplish this, we propose a cross-modal pipeline that generates high-quality cognitive video CoT data simulating human reasoning processes, resulting in two large-scale datasets: VideoRFT-CoT-102K and VideoRFT-RL-310K.
+Furthermore, to strengthen the RL phase, we develop semantic-consistency guided reward to explicitly encourage the alignment between reasoning traces and visual evidence.
+Extensive experiments across six benchmarks demonstrate that VideoRFT consistently surpasses a variety of advanced MLLMs.
 We expect this work to lay a foundation for future efforts in RFT-based video reasoning.